Determine whether or not the given map is an isomorphism on the structures . If it isn’t explain why. Let F be the set of all functions f mapping R —> R that have derivatives of all orders . <F,+> with <R,+) with phi (f) = f’(0) for f is an element of F


1
Expert's answer
2021-02-12T18:09:11-0500

Consider the structures "\langle F,+\rangle" and "\langle \mathbb R,+\rangle", and the map "\psi: F\to \mathbb R""\psi (f) = f'(0)". This map is not injective. Indeed, for the different functions "f_1(x)=2e^x" and "f_2(x)=e^{2x}" that have derivatives of all orders, in particular, "f_1'(x)=2e^x" and "f_2'(x)=2e^{2x}", we have that "\psi (f_1) = f_1'(0)=2e^0=2" and "\psi (f_2) = f_2'(0)=2e^{2\cdot 0}=2." Therefore, "\psi (f_1) =\psi (f_2)", and "\psi" is not a injection. Consequently, "\psi" is not an isomorphism.